# FLEX: Continuous Agent Evolution via Forward Learning from Experience

### v1

@@ 6.2 Self-Evolving Agent
Recent research has begun to explore the vision of self-evolving agents [gao2025survey, fang2025comprehensive], that can autonomously improve their reasoning and decision-making through continual interaction with the environment. A first line of work focuses on tool evolution, where agents autonomously master, create, and utilize tools to extend their capabilities, such as Toolformer [schick2023toolformer], ReAct [yao2022react], Voyager [wang2023voyager], CREATOR [qian2023creator], and ALITA [qiu2025alita]. Another line emphasizes agents architecture evolution, such as CAMEL [li2023camel], MetaGPT [hong2023metagpt], AgentSquare [shang2024agentsquare], MaAS [zhang2025multi], AutoFlow [li2024autoflow], and GPTSwarm [zhuge2024gptswarm], which orchestra multiple agents for cooperative problem solving. A third line focuses on context evolution, where agents continually refine their internal reasoning via reflection and semantic feedback Representative studies include Reflexion [shinn2023reflexion], Self-Refine [madaan2023self], GEPA [agrawal2025gepa], SE-Agent [lin2025se], and ACE [zhang2025agentic]. Similarly, TextGrad [yuksekgonul2024textgrad] and REVOLVE [zhang2024revolve] interpret semantic feedback as textual gradients to guide model refinement.
Recently, experience-driven evolution [tang2025agent, zhou2025memento, ouyang2025reasoningbank, cai2025training] has emerged, allowing LLM agents to accumulate and reuse past interaction trajectories. AgentKB [tang2025agent] emphasizes cross-domain knowledge transfer through a knowledge base. Memento [zhou2025memento] stores raw reasoning trajectories for later retrieval, but lacks generalization and interpretive guidance for novel contexts. ReasoningBank [ouyang2025reasoningbank] scales memory retrieval at inference time without introducing an explicit learning process. TF-GRPO [cai2025training] simply mimics the GRPO [shao2024deepseekmath] procedure to compute group-relative semantic advantages, but its extensibility and generalization are constrained by the GRPO-style design.
@@ 7 Conclusion
[…]
8 Experimental Details
8.1 Biology
Protein Fitness Prediction We apply FLEX on the task of Protein Fitness Prediction. Proteins play a fundamental role in sustaining cellular functions that underpin organismal survival, growth, and reproduction. The capacity of a protein to perform its biological function—typically referred to as protein fitness—is determined by its three-dimensional structure, which is ultimately encoded by its amino acid sequence. Recent advances in machine learning have enabled zero-shot protein fitness prediction by leveraging representations derived from protein sequences, structural information, and auxiliary biological signals such as multiple sequence alignments (MSAs). Despite the progress, the task remains far from being fully solved. First, the absolute correlation between predicted and experimentally validated fitness scores remains limited (e.g., approximately 0.52 for the current state-of-the-art on ProteinGym), not confidential enough when encountering newly-found data with mysterious fitness. Second, no existing model consistently outperforms all others across diverse protein families and functional categories, including activity, expression, binding affinity, stability, and organismal fitness.
Agentic Task Transformation.
Large language model (LLM) agents, without specialized pre-training on protein sequences or domain-specific fine-tuning, inherently lack detailed biochemical knowledge and cannot directly infer subtle mutational effects on protein function. However, these agents possess strong capabilities in reasoning, ranking, knowledge synthesis, and few-shot learning. This opens a new opportunity: rather than directly predicting fitness from sequence, agents can be tasked with meta-reasoning over predictions made by high-performing protein language models. Specifically, we reformulate protein fitness prediction as an agentic feature selection and fusion task, where the agent learns to identify, prioritize, and integrate salient features derived from existing regression-based models. Through this process, the agent proposes explicit, interpretable algorithms that capture the underlying relationships among predictive signals, transforming opaque model outputs into a more accurate and explainable fitness prediction framework tailored to each target protein.
Benchmark.
We conduct experiments using the ProteinGym benchmark, which serves both as the source for experience extraction and as the evaluation dataset. Protein targets included in this benchmark align with those defined in SaProt [su2023saprot]. To emulate realistic biological discovery scenarios where experimental data is scarce, we adopt an imbalanced data split for each protein: a small experience set consisting of only 100 sequences (on average 1.47% of the available data), and a large out-of-distribution test set containing all remaining sequences. This design reflects practical constraints in protein engineering, where only a limited number of labeled mutations are known, while the vast majority of sequence space remains unexplored.
Figure 7: ProteinGym Results from both protein language model and large language model.
Setup.
Prediction performance is evaluated using the Spearman rank correlation between the ground truth DMS (deep mutational scanning) scores and the proxy metrics generated by our FLEX method. Baseline models are grouped into two main categories:
Zero-shot Protein Fitness Prediction Models. These models represent the current state-of-the-art on the ProteinGym leader-board and provide diverse biological priors through different input modalities:
  • VespaG [marquet2024expert, marquet2022embeddings]: A top-performing single-sequence model leveraging PLM embeddings and evolutionary guidance from GEMME [laine2019gemme].
  • PoET [truong2023poet]: A homolog-based model that encodes multiple sequence alignments using a sequence-of-sequences transformer architecture.
  • ProSST [li2024prosst]: A hybrid model incorporating structural quantization and disentangled attention mechanisms to jointly model sequence and structural representations.
  • VenusREM [tan2025high]: The current state-of-the-art on the ProteinGym scoreboard, integrating logits from single-sequence, homology, and structural tokenization signals.
These models provide the input features that our agent uses to reason about and construct enhanced fitness metrics.
Native Large Language Models (LLMs). To isolate the contribution of agentic reasoning, we evaluate several high-performance LLMs—such as GPT-OSS [agarwal2025gpt] and Claude-Sonnet-4.5 as direct baselines. These models are prompted to perform the same prediction task as our forward learning system but without employing experience validation, rejection sampling and refinement. This comparison highlights the added value of structured forward learning beyond raw language model capabilities.
Table 3: Ablation study results in ProteinGym benchmark of forward learning paradigm component techniques: Experience Exploration, Experience Evolution, Assistance of Regression tools. Where Fixed mean best one for all regression tools is used instead of each different regression tools for protein target selected by agent.
Model Experience Exploration Experience Evolution Regression Tools Spearman’s $\rho$
GPT-OSS 0.472
Fixed 0.531
0.537
0.547
0.573
Fixed 0.568
0.581
Ablation Study. From Table 3 we can learn that, three core parts of our FLEX paradigm contributed to performance enhancement: Experience Exploration, Experience Evolution, Regression Tools. Specificlly, Experience Exploration provide success and failure cases agent used to engage to, in the ProteinGym Scenario, common features included in most successful cases suggest its effect-ness in prediction final proxy metric. Experience Evolution enables agent to summarize past experience, update lessons learned recently with new cross-validation facts. Moreover, based on accurate selection of top importance features, regression tools could help agents to further adjust limited hyper-parameters, where deciding what kinds of regression works best still important for agents compared to the fix, one for all regression cases.

### v2

@@ 6.2 Self-Evolving Agent
Recent research has begun to explore the vision of self-evolving agents [110, 111], that can autonomously improve their reasoning and decision-making through continual interaction with the environment. A first line of work focuses on tool evolution, where agents autonomously master, create, and utilize tools to extend their capabilities, such as Toolformer [112], ReAct [38], Voyager [113], CREATOR [27], and ALITA [28]. Another line emphasizes agents architecture evolution, such as CAMEL [114], MetaGPT [115], AgentSquare [26], MaAS [25], AutoFlow [116], and GPTSwarm [117], which orchestra multiple agents for cooperative problem solving. A third line focuses on context evolution, where agents continually refine their internal reasoning via reflection and semantic feedback Representative studies include Reflexion [29], Self-Refine [37], GEPA [118], SE-Agent [24], and ACE [23]. Similarly, TextGrad [22] and REVOLVE [30] interpret semantic feedback as textual gradients to guide model refinement.
Recently, experience-driven evolution [119, 120, 121, 14] has emerged, allowing LLM agents to accumulate and reuse past interaction trajectories. AgentKB [119] emphasizes cross-domain knowledge transfer through a knowledge base. Memento [120] stores raw reasoning trajectories for later retrieval, but lacks generalization and interpretive guidance for novel contexts. ReasoningBank [121] scales memory retrieval at inference time without introducing an explicit learning process. TF-GRPO [14] simply mimics the GRPO [91] procedure to compute group-relative semantic advantages, but its extensibility and generalization are constrained by the GRPO-style design.
@@ 7 Conclusion
Acknowledgments and Disclosure of Funding
This work is supported by the National Key R&D Program of China (2022ZD0160501).